OpenStudy (tester97):

Rewrite cubic root(27x-81) -5 to make it easier to graph using the translation.

OpenStudy (tester97):

\[\sqrt[3]{27x-81}-5\]

OpenStudy (kc_kennylau):

Take out 27 from 27x-81 :)

OpenStudy (kc_kennylau):

\[\sqrt[3]{27x-81}-5=\sqrt[3]{27(x-3)}-5=\sqrt[3]{27}\sqrt[3]{x-3}-5=3\sqrt[3]{x-3}-5\]

OpenStudy (tester97):

thanks kc :)

OpenStudy (anonymous):

hi directrix

OpenStudy (kc_kennylau):

Welcome to OpenStudy @trollwillrule :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (kc_kennylau):

@tester97 no problem :)